# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v2

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
see [42, Scetion 5]. Using (4.8) and (4.14), we may view every factor in (4.13) as a function of the same self-adjoint operator $\sqrt{-\mathcal{L}}$. In spectral variable $\zeta\geq 0$ (corresponding to $\sqrt{-\mathcal{L}}$), we obtain
@@
Let $\mathcal{S}_{\sigma}^{m}$ denote the standard strip-holomorphic symbol class appearing in the $L^{p}$ multiplier theorem on noncompact manifolds with bounded geometry: namely $\Phi\in\mathcal{S}_{\sigma}^{m}$ if $\Phi$ is an even function and extends holomorphically to $\{\zeta\in\mathbb{C}:|{\rm Im}(\zeta)|<\sigma\}$ and satisfies $|\partial_{\zeta}^{k}\Phi(\zeta)|\leq C_{k}(1+|\zeta|)^{m-k}$ there. By [42, Chapter IV, Proposition 2.4], it is enough to show that there exists $\sigma>0$ such that
@@
Hence $m_{\gamma}\in\mathcal{S}_{\frac{\epsilon}{2}}^{2\gamma}$. Multiplying the two symbols yields (4.15), see [20, Chapter 18]
@@
By [42, Chapter IV, Proposition 1.2], it is equivalent to $\mathcal{P}_{s}\in\Psi_{\sigma}^{2s}(\mathbb{H}^{n}),$ where the space $\Psi_{W}^{m}(\mathbb{H}^{n})$, of a class of operators whose Schwartz kernels behave like those of operators in $\mathcal{S}^{m}_{0,1}$ near the diagonal, in a uniform fashion, and away from the diagonal decay like $d(x,y)^{-k}e^{-Wd(x,y)},\ \forall k$ as do all derivatives. See [42, Page 72].
@@
Finally, we use the Sobolev mapping property for pseudo-differential operators ([42, Proposition 4.5]): for ${\bf T}\in\Psi_{\sigma}^{m}(\mathbb{H}^{n})$ and any $t\in\mathbb{R}$,
@@
    Let $s\in\bigl(0,\frac{n}{2}\bigr)\setminus\mathbb{N}$ and $\lambda<\lambda_{0,s}^{\mathrm{conf}}$. Then
@@
    Let $s\in\bigl(0,\frac{n}{2}\bigr)$ and $\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$. Then
@@
Let $\mathcal{A}_{\lambda}:=(\lambda I-\Delta_{\mathbb{H}^{n}})^{1/2}$, so that $\|u\|_{\widetilde{H}^{t}}=\|\mathcal{A}_{\lambda}^{t}u\|_{L^{2}}$ for all $t\in\mathbb{R}$. Let $\{v_{j}\}\subset\widetilde{H}^{s}(\mathbb{H}^{n})$ be bounded and set $f_{j}:=\mathcal{A}_{\lambda}^{s}v_{j}\in L^{2}(\mathbb{H}^{n})$. Then
$\|\psi v_{j}\|_{\widetilde{H}^{s-1}}=\|\mathcal{A}_{\lambda}^{s-1}(\psi v_{j})\|_{L^{2}}=\|{\bf K}_{\psi}f_{j}\|_{L^{2}},$
@@
is compact. Since $v_{j}\rightharpoonup 0$ weakly in $\tilde{H}^{s}(\mathbb{H}^{n})=\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, $f_{j}\rightharpoonup 0$ weakly in $L^{2}(\mathbb{H}^{n})$ as $j\to+\infty$, it follows that ${\bf K}_{\psi}f_{j}\to 0$ strongly in $L^{2}$. Hence
$\|\psi v_{j}\|_{\tilde{H}^{s-1}(\mathbb{H}^{n})}\longrightarrow 0\quad{\rm as}\ \,j\to+\infty.$
@@
Proof of Theorem 1.8. Since $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, we have $H_{n,s,p}(\lambda)>0$ by (4.6). Therefore, it suffices to establish the existence of minimizers for $H_{n,s,p}(\lambda)$.
@@
$I_{\lambda,s}(u_{j})\to H_{n,s}(\lambda)\quad{\rm as}\ j\to+\infty.$
@@
$\|u_{j}\|_{p+1}^{p+1}=\mathcal{E}_{\lambda,s}(u_{j})=I_{\lambda,s}(u_{j})^{\frac{p+1}{p-1}}\to H_{n,s}(\lambda)^{\frac{p+1}{p-1}}.$ (4.18)
@@
$0<\delta<H_{n,s}(\lambda)^{\frac{p+1}{p-1}}.$
@@
$\liminf_{j\to\infty}\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\geq H_{n,s}(\lambda)^{\frac{p+1}{p-1}}.$ (4.26)
Recall that in (4.20) we fixed $0<\delta<H_{n,s}(\lambda)^{\frac{p+1}{p-1}}$, this contradicts (4.26). Therefore necessarily $\|\phi v_{j}\|_{p+1}\to 0$, which is exactly (4.22).
@@
We use the Harish–Chandra expansion from [40, Theorem 3.1] (for rank one) which, in our notation, reads
@@
For each fixed $r\geq 1$, the series converges uniformly in $\beta$ on compact sets [40, Remark 1 after Theorem 3.1]. Moreover, the coefficients $\Gamma_{k}(\beta)$ have at most polynomial growth in $\beta$ (see the discussion following Theorem 3.1 and the estimates on $\Gamma_{k}$ in Section 3 of [40]), hence the factor $e^{-2r}$ implies that $J_{+}^{(1)}(r)$ enjoys the same oscillatory integration–by–parts bound as $J_{+}^{(0)}(r)$, up to an additional $e^{-2r}$ factor. Concretely, repeating the above argument with the amplitude $a(\beta)\big(\phi_{\beta}(r)-1\big)$ yields
@@
Proof of Theorem 1.5: (i) By Theorem 1.3 (i) and Proposition 4.12, we obtain $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$. By the proof of Theorem 1.7, $\widetilde{H}_{n,s}(\lambda)$ is achieved for every $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$, thus by Lemma 5.5 (ii), $\widetilde{H}_{n,s}(\lambda)$ is strictly decreasing in $(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$. The remaining conclusions follow from Proposition 4.11, Proposition 4.12 and Lemma 5.4.
@@
(iii) By [29, Theorem 1.9], we obtain $\widetilde{H}_{n,s}\!\bigl(\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}.$ By Proposition 4.12, $\widetilde{H}_{n,s}\!\bigl(0\bigr)=S_{n,s},$ thus we complete the proof. $\Box$
@@
(ii) For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, we have $\lambda_{0,s}^{\mathrm{conf}}>b_{s}>0$. By Proposition 4.12 and (1.6) we know $S_{n,s}=\widetilde{H}_{n,s}(0)\leq H_{n,s}(0)\leq S_{n,s},$ thus $H_{n,s}(0)=S_{n,s}.$ Then by Theorem 1.4 (i), we have $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr].$ By the proof of Theorem 1.8 and Lemma 5.5 (ii), $H_{n,s}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}})$.
@@
(iv) By [29, Theorem 1.4] and Proposition 4.12, for $\lambda\leq{\lambda}_{0,s}^{\mathrm{conf}}$, we obtain $H_{n,s}\!\bigl({\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}=H_{n,s}(\lambda).$ The remaining conclusions follow from Proposition 4.11. $\Box$
@@ 5.1 Stability and Attainability of Sobolev Levels
(i) Using [33, Claim 14.1], Proposition 5.2, Lemma 5.5, the proof is completely analogous to the case $s=1$, and we therefore omit the details.
(ii) By [33, Proposition 16.4], we obtain $(\lambda_{s}^{*},\infty)\subset\mathcal{G}_{n,s}[S_{n,s}]$ and thus, $S_{n,s}(\lambda)$ is attained if $\lambda\in(\lambda_{s}^{*},\infty)$. By Proposition 5.2, Lemma 5.5, for any $\lambda\leq 0<\lambda_{s}^{*}<\mu_{1}<\mu_{2}$,
@@
By the proof of [30, Theorem 1.5, Theorem 1.6], we know that $H_{n,1}(\lambda)$ is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}].$ Thus, by Lemma 5.5(ii), for any $\lambda\leq 0<\mu_{1}<\mu_{2}<\frac{1}{4},$ we have
@@
Thus, we prove the inequality and $\mathcal{G}_{3,1}[H_{3,1}]=(\frac{1}{4},\infty).$ By [30, Theorem 1.7], we obtain $H_{3,1}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
@@
(i) By [25, Theorem 1.14] and [28, Theorem 1.7], we know that $H_{n,k}(\lambda)<S_{n,k}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $n\geq 4k$. By arguments completely analogous to those used in the preceding propositions in high dimensions, we obtain the desired conclusion here, and therefore omit the details.
(ii) By [25, Theorem 1.14] and [28, Theorem 1.7], there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $H_{n,k}(\lambda)<S_{n,k}$ when $\lambda>\lambda_{k}^{\mathrm{conf}}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$. The remainder of the proof follows by an entirely analogous argument, combined with Lemma 5.5 (ii) and Proposition 4.11, and we therefore omit the details.
(iii) When $n=2k+1$, by [29, Theorem 1.6] or [28, Theorem 1.2], we obtain $H_{n,k}(\lambda_{0,k}^{\mathrm{conf}})=S_{n,k}.$ The rest of the proof is completely analogous to (ii), and we omit it. $\Box$

### v3

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
see [43, Scetion 5]. Using (4.8) and (4.14), we may view every factor in (4.13) as a function of the same self-adjoint operator $\sqrt{-\mathcal{L}}$. In spectral variable $\zeta\geq 0$ (corresponding to $\sqrt{-\mathcal{L}}$), we obtain
@@
Let $\mathcal{S}_{\sigma}^{m}$ denote the standard strip-holomorphic symbol class appearing in the $L^{p}$ multiplier theorem on noncompact manifolds with bounded geometry: namely $\Phi\in\mathcal{S}_{\sigma}^{m}$ if $\Phi$ is an even function and extends holomorphically to $\{\zeta\in\mathbb{C}:|{\rm Im}(\zeta)|<\sigma\}$ and satisfies $|\partial_{\zeta}^{k}\Phi(\zeta)|\leq C_{k}(1+|\zeta|)^{m-k}$ there. By [43, Chapter IV, Proposition 2.4], it is enough to show that there exists $\sigma>0$ such that
@@
Hence $m_{\gamma}\in\mathcal{S}_{\frac{\epsilon}{2}}^{2\gamma}$. Multiplying the two symbols yields (4.15), see [21, Chapter 18]
@@
By [43, Chapter IV, Proposition 1.2], it is equivalent to $\mathcal{P}_{s}\in\Psi_{\sigma}^{2s}(\mathbb{H}^{n}),$ where the space $\Psi_{W}^{m}(\mathbb{H}^{n})$, of a class of operators whose Schwartz kernels behave like those of operators in $\mathcal{S}^{m}_{0,1}$ near the diagonal, in a uniform fashion, and away from the diagonal decay like $d(x,y)^{-k}e^{-Wd(x,y)},\ \forall k$ as do all derivatives. See [43, Page 72].
@@
Finally, we use the Sobolev mapping property for pseudo-differential operators ([43, Proposition 4.5]): for ${\bf T}\in\Psi_{\sigma}^{m}(\mathbb{H}^{n})$ and any $t\in\mathbb{R}$,
@@
    Let $s\in\bigl(0,\frac{n}{2}\bigr)\setminus\mathbb{N}$ and $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$. Then
@@
    Let $s\in\bigl(0,\frac{n}{2}\bigr)$ and $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$. Then
@@
Choose $\eta\in C_{c}^{\infty}(\mathbb{H}^{n})$ such that
$\eta\equiv 1\quad\text{on a neighborhood of }\operatorname{supp}\psi,$
and define $w_{j}:=\eta v_{j}.$ Then $\operatorname{supp}w_{j}\subset\operatorname{supp}\eta=:K\Subset\mathbb{H}^{n}$ for all $j$, so $\{w_{j}\}$ has a common compact support. Moreover,
$\psi w_{j}=\psi(\eta v_{j})=(\psi\eta)v_{j}=\psi v_{j},$
hence
$\|\psi v_{j}\|_{\tilde{H}^{\,s-1}(\mathbb{H}^{n})}=\|\psi w_{j}\|_{\tilde{H}^{\,s-1}(\mathbb{H}^{n})}.$
Therefore, the commutator estimate becomes
$\|[\mathcal{P}_{s},\psi](\psi v_{j})\|_{\tilde{H}^{-s}(\mathbb{H}^{n})}\leq C_{\psi}\,\|\psi v_{j}\|_{\tilde{H}^{\,s-1}(\mathbb{H}^{n})}=C_{\psi}\,\|\psi w_{j}\|_{\tilde{H}^{\,s-1}(\mathbb{H}^{n})}.$
Since $v_{j}\rightharpoonup 0$ weakly in the critical energy space and multiplication by a fixed cutoff, we also have
$w_{j}=\eta v_{j}\rightharpoonup 0\quad\text{weakly in }\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}).$
By Remark 4.1 and Lemma 4.4 (ii), it follows that for $\lambda\leq\lambda_{0,s}^{\mathrm{conf}},$
$w_{j}\rightharpoonup 0\quad\text{weakly in }H^{s}(\mathbb{H}^{n}).$ (4.18)
Let $\mathcal{A}_{\lambda}:=(\lambda I-\Delta_{\mathbb{H}^{n}})^{1/2}$, so that $\|u\|_{\widetilde{H}^{t}}=\|\mathcal{A}_{\lambda}^{t}u\|_{L^{2}}$ for all $t\in\mathbb{R}$. Let $\{w_{j}\}\subset\widetilde{H}^{s}(\mathbb{H}^{n})$ be bounded and set $f_{j}:=\mathcal{A}_{\lambda}^{s}w_{j}\in L^{2}(\mathbb{H}^{n})$. Then
$\|\psi w_{j}\|_{\widetilde{H}^{s-1}}=\|\mathcal{A}_{\lambda}^{s-1}(\psi w_{j})\|_{L^{2}}=\|{\bf K}_{\psi}f_{j}\|_{L^{2}},$
@@
is compact. By (4.18), $f_{j}\rightharpoonup 0$ weakly in $L^{2}(\mathbb{H}^{n})$ as $j\to+\infty$, it follows that ${\bf K}_{\psi}f_{j}\to 0$ strongly in $L^{2}$. Hence
$\|\psi w_{j}\|_{\tilde{H}^{s-1}(\mathbb{H}^{n})}\longrightarrow 0\quad{\rm as}\ \,j\to+\infty.$
@@
Proof of Theorem 1.8. Since $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$, we have $H_{n,s,p}(\lambda)>0$ by (4.6). Therefore, it suffices to establish the existence of minimizers for $H_{n,s,p}(\lambda)$.
@@
$I_{\lambda,s}(u_{j})\to H_{n,s,p}(\lambda)\quad{\rm as}\ j\to+\infty.$
@@
$\|u_{j}\|_{p+1}^{p+1}=\mathcal{E}_{\lambda,s}(u_{j})=I_{\lambda,s}(u_{j})^{\frac{p+1}{p-1}}\to H_{n,s,p}(\lambda)^{\frac{p+1}{p-1}}.$ (4.19)
@@
$0<\delta<H_{n,s,p}(\lambda)^{\frac{p+1}{p-1}}.$
@@
$\liminf_{j\to\infty}\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\geq H_{n,s,p}(\lambda)^{\frac{p+1}{p-1}}.$ (4.27)
Recall that in (4.21) we fixed $0<\delta<H_{n,s,p}(\lambda)^{\frac{p+1}{p-1}}$, this contradicts (4.27). Therefore necessarily $\|\phi v_{j}\|_{p+1}\to 0$, which is exactly (4.23).
@@
We use the Harish–Chandra expansion from [41, Theorem 3.1] (for rank one) which, in our notation, reads
@@
For each fixed $r\geq 1$, the series converges uniformly in $\beta$ on compact sets [41, Remark 1 after Theorem 3.1]. Moreover, the coefficients $\Gamma_{k}(\beta)$ have at most polynomial growth in $\beta$ (see the discussion following Theorem 3.1 and the estimates on $\Gamma_{k}$ in Section 3 of [41]), hence the factor $e^{-2r}$ implies that $J_{+}^{(1)}(r)$ enjoys the same oscillatory integration–by–parts bound as $J_{+}^{(0)}(r)$, up to an additional $e^{-2r}$ factor. Concretely, repeating the above argument with the amplitude $a(\beta)\big(\phi_{\beta}(r)-1\big)$ yields
@@
Proof of Theorem 1.5: (i) By Theorem 1.3 (i) and Proposition 4.13, we obtain $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$. By the proof of Theorem 1.7, $\widetilde{H}_{n,s}(\lambda)$ is achieved for every $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$, thus by Lemma 5.5 (ii), $\widetilde{H}_{n,s}(\lambda)$ is strictly decreasing in $(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$. The remaining conclusions follow from Proposition 4.12, Proposition 4.13 and Lemma 5.4.
@@
(iii) By [30, Theorem 1.9], we obtain $\widetilde{H}_{n,s}\!\bigl(\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}.$ By Proposition 4.13, $\widetilde{H}_{n,s}\!\bigl(0\bigr)=S_{n,s},$ thus we complete the proof. $\Box$
@@
(ii) For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, we have $\lambda_{0,s}^{\mathrm{conf}}>b_{s}>0$. By Proposition 4.13 and (1.6) we know $S_{n,s}=\widetilde{H}_{n,s}(0)\leq H_{n,s}(0)\leq S_{n,s},$ thus $H_{n,s}(0)=S_{n,s}.$ Then by Theorem 1.4 (i), we have $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr].$ By the proof of Theorem 1.8 and Lemma 5.5 (ii), $H_{n,s}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}}]$.
@@
(iv) By [30, Theorem 1.4] and Proposition 4.13, for $\lambda\leq{\lambda}_{0,s}^{\mathrm{conf}}$, we obtain $H_{n,s}\!\bigl({\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}=H_{n,s}(\lambda).$ The remaining conclusions follow from Proposition 4.12. $\Box$
@@ 5.1 Stability and Attainability of Sobolev Levels
(i) Using [34, Claim 14.1], Proposition 5.2, Lemma 5.5, the proof is completely analogous to the case $s=1$, and we therefore omit the details.
(ii) By [34, Proposition 16.4], we obtain $(\lambda_{s}^{*},\infty)\subset\mathcal{G}_{n,s}[S_{n,s}]$ and thus, $S_{n,s}(\lambda)$ is attained if $\lambda\in(\lambda_{s}^{*},\infty)$. By Proposition 5.2, Lemma 5.5, for any $\lambda\leq 0<\lambda_{s}^{*}<\mu_{1}<\mu_{2}$,
@@
By the proof of [31, Theorem 1.5, Theorem 1.6], we know that $H_{n,1}(\lambda)$ is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}].$ Thus, by Lemma 5.5(ii), for any $\lambda\leq 0<\mu_{1}<\mu_{2}<\frac{1}{4},$ we have
@@
Thus, we prove the inequality and $\mathcal{G}_{3,1}[H_{3,1}]=(\frac{1}{4},\infty).$ By [31, Theorem 1.7], we obtain $H_{3,1}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
@@
(i) By [26, Theorem 1.14] and [29, Theorem 1.7], we know that $H_{n,k}(\lambda)<S_{n,k}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $n\geq 4k$. By arguments completely analogous to those used in the preceding propositions in high dimensions, we obtain the desired conclusion here, and therefore omit the details.
(ii) By [26, Theorem 1.14] and [29, Theorem 1.7], there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $H_{n,k}(\lambda)<S_{n,k}$ when $\lambda>\lambda_{k}^{\mathrm{conf}}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$. The remainder of the proof follows by an entirely analogous argument, combined with Lemma 5.5 (ii) and Proposition 4.12, and we therefore omit the details.
(iii) When $n=2k+1$, by [30, Theorem 1.6] or [29, Theorem 1.2], we obtain $H_{n,k}(\lambda_{0,k}^{\mathrm{conf}})=S_{n,k}.$ The rest of the proof is completely analogous to (ii), and we omit it. $\Box$